# Granite-speech: open-source speech-aware LLMs with strong English ASR capabilities

### v1

@@ I Introduction
Our proposed speech-aware LLMs operate in two steps. In step one, English speech input is transcribed or, optionally, translated into a different language. The output of step one can be fed, in step two, as a text prompt, potentially with a longer dialog context, to the underlying Granite text LLM to generate a response. Importantly, the text LLM is shared between the two steps, requiring only one instance in memory. During inference, we support two modes of operation: a text-only mode and a speech mode. The speech mode is automatically activated when a prompt includes both an ¡—audio—¿ token and a corresponding audio file; otherwise, the model defaults to text-only mode. This setup allows users to toggle modes seamlessly by simply including or omitting the audio input. In speech mode, the model runs with the acoustic encoder, speech modality adapter, and enables the LoRA adapter. In text-only mode, the text LLM is used with LoRA adapters disabled. By construction, this two-pass architecture ensures that all text capabilities such as retrieval-augmented generation, function-calling, and safety are preserved, at the cost of requiring two LLM generation calls. This disadvantage can be mitigated through careful orchestration of the calls, for example, by caching previous key-value computations in the LLM attention layers. Compared to a cascaded dedicated ASR+LLM approach, our proposed model has the advantage of a stronger ASR component by leveraging the power of the LLM. We position the granite-speech-3.3 models as LLMs with good ASR/AST capabilities rather than standalone ASR/AST models (although they can certainly be used this way).
@@ II-D Task-specific prompt construction
The user query can correspond to either a transcription (ASR) or a translation (AST). Each training example is labeled with a task tag, which determines how the prompt will be selected. Prompts contain a special ¡—audio—¿ token, which gets replaced with the projected embeddings of the input audio. Given an ASR example, we randomly select a prompt from a set of 24 variations, e.g.:
Listen to the speech and write down its content ¡—audio—¿.
@@
¡—audio—¿ translate the speech to Spanish.
@@
¡—audio—¿ Can you transcribe the speech, and then translate it to Spanish?
@@
After constructing the final textual input, we tokenize and embed it using the Granite tokenizer and text embedding table. We then replace the ¡—audio—¿ special token with the projected embeddings of the audio. This combined representation is then passed to the LLM for generating the corresponding output. To enable the LLM to consume a compact and informative representation of the input audio, we next describe the architecture of our modality adapter.

### v2

@@ I Introduction
Our proposed speech-aware LLMs operate in two steps. In step one, English speech input is transcribed or, optionally, translated into a different language. The output of step one can be fed, in step two, as a text prompt, potentially with a longer dialog context, to the underlying Granite text LLM to generate a response. Importantly, the text LLM is shared between the two steps, requiring only one instance in memory. During inference, we support two modes of operation: a text-only mode and a speech mode. The speech mode is automatically activated when a prompt includes both an <|audio|> token and a corresponding audio file; otherwise, the model defaults to text-only mode. This setup allows users to toggle modes seamlessly by simply including or omitting the audio input. In speech mode, the model runs with the acoustic encoder, speech modality adapter, and enables the LoRA adapter. In text-only mode, the text LLM is used with LoRA adapters disabled. By construction, this two-pass architecture ensures that all text capabilities such as retrieval-augmented generation, function-calling, and safety are preserved, at the cost of requiring two LLM generation calls. This disadvantage can be mitigated through careful orchestration of the calls, for example, by caching previous key-value computations in the LLM attention layers. Compared to a cascaded dedicated ASR+LLM approach, our proposed model has the advantage of a stronger ASR component by leveraging the power of the LLM. We position the granite-speech-3.3 models as LLMs with good ASR/AST capabilities rather than standalone ASR/AST models (although they can certainly be used this way).
@@ II-D Task-specific prompt construction
The user query can correspond to either a transcription (ASR) or a translation (AST). Each training example is labeled with a task tag, which determines how the prompt will be selected. Prompts contain a special <|audio|> token, which gets replaced with the projected embeddings of the input audio. Given an ASR example, we randomly select a prompt from a set of 24 variations, e.g.:
Listen to the speech and write down its content <|audio|>.
@@
<|audio|> translate the speech to Spanish.
@@
<|audio|> Can you transcribe the speech, and then translate it to Spanish?
@@
After constructing the final textual input, we tokenize and embed it using the Granite tokenizer and text embedding table. We then replace the <|audio|> special token with the projected embeddings of the audio. This combined representation is then passed to the LLM for generating the corresponding output. To enable the LLM to consume a compact and informative representation of the input audio, we next describe the architecture of our modality adapter.